OpenStudy (anonymous):

A grocer bought a box of soap containing 72 bars for $49.00. Some of the soap he sold at 3 bars for $2.50, and the rest $0.70 a bar. This gave a profit of $7.80. How many bars did he sell?

OpenStudy (anonymous):

Anybody wanna help pleas???

OpenStudy (anonymous):

i'll try =^_^=

OpenStudy (anonymous):

hold on..

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

$2.5 = 48 bars $0.7 = 24 bars

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

happy to help! =^_^=

OpenStudy (anonymous):

it is not ..

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

sure pat?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

yeah, i got that also. sorry.

OpenStudy (anonymous):

haha..

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

Let x = number sold at 3 / $2.50 72 - x = number sold at $.70 2.50 / 3x + 2.10 (72- x = 49.00 + 7.80 2.50x + 151.20 - 2.10x = 170.40 .4x = 19.20 x = 48 48 \[\div\] 3 = 16 2.50 x 16 = 40 .70 x 24 = 16.80 40.00 + 16.80 = 56.80 56.80 - 49.00 = 7.80

OpenStudy (anonymous):

Let x=number sold at 0.7 72-x=number sold at 0.25 2.5(72-x)/3 + 0.7x=56.8 180-2.5x+2.1x=170.4 -0.4x=-9.6 x=24 72-x=72-24=48